OpenStudy (anonymous):

Classify the model as exponential growth or decay. Identify growth or decay factor and the percent of increase or decrease per time period. a) y=40(0.75)^t please explain

OpenStudy (anonymous):

since \(.75<1\) it is decay

OpenStudy (anonymous):

multiplying a number by \(.75\) decreases it by 25%

OpenStudy (anonymous):

and how do you find the second part?

OpenStudy (anonymous):

the second post was the second part

OpenStudy (anonymous):

what if it is 6/5 instead of .75, how would you find that?

OpenStudy (agent0smith):

6/5 is greater than 1, so it's exponential growth. The growth factor will be\[ 6/5 - 1 = 1.2 - 1 = ? \]

OpenStudy (anonymous):

.2

OpenStudy (agent0smith):

Sorry, the growth factor will be the 6/5 or 1.2. The percent of increase per time period will be \[(1.2−1)×100 % \]

OpenStudy (anonymous):

ok thanks alot